OpenStudy (anonymous):

Divide x2 + 3x - 18 by x - 3

OpenStudy (anonymous):

x= -6

OpenStudy (solomonzelman):

no, it's not solving, it's dividing.

OpenStudy (anonymous):

x^2 + 3x - 18/(x-3)=(x+6)(x-3)/(x-3)=x+6

OpenStudy (anonymous):

mustafa2014 is correct